An ice hockey rink is a rectangle with a perimeter of 174 meters. The length is 9 meters more than twice the width. Find the length and width?
Found 2 solutions by stanbon, Adrish:
Answer by stanbon(75887)

An ice hockey rink is a rectangle with a perimeter of 174 meters. The length is 9 meters more than twice the width. Find the length and width?
---
Perimeter = 2(width + length)
Let width = x
Then length = 2x+9
--------
174 = 2(x + 2x+9)
87 = 3x+9
3x = 78
x = 26
----
Width = 26 meters
Length = 2x+9 = 52+9 = 61 meters
----------------------------
Cheers,
Stan H.
=========

Answer by Adrish(44)

Solutions: Let the width be "x"
and the length be 9+2w
Given:Perimeter=174 meters
we know Perimeter of rectangle=2L + 2W
P=2Lx 2W
174=2(9+2w) + 2w
174=18 +4w +2w
174=18+6w
6w=174-18
6w=156
w=156/6
w=26
Therefore Width=26 meter
Length=9+2w=9+2x26
=9+ 52
=61 meter